OpenStudy (anonymous):

A right triangle has a short side of 15m and bottom side of 20m. How do I find perimeter?

OpenStudy (anonymous):

assuming the "bottom side" isn't the hypotenuse (long side), use Pythagorean theorum a^2+b^2=c^2 225+400=c^2 25=c 15+20+25=perimeter=60

OpenStudy (anonymous):

Thank you, thank you very much........